# Deconfined quantum critical points in fermionic systems with spin-charge separation

### v3

@@ Abstract
Deconfined quantum critical points are intriguing transition points not predicted by the Landau–Ginzburg-Wilson symmetry-breaking paradigm which are usually identified by the appearance of a continuous phase transitions between locally ordered phases. Here, we reveal the presence of deconfined quantum critical points with unexplored properties. Contrary to previously known examples, we indeed show that the phenomenon of spin-charge separation peculiar of interacting low dimensional fermions can allow for the appearance of partially gapped deconfined quantum critical points. We first infer this point by performing a field theory analysis of generic one-dimensional fermionic systems in the low energy limit. Subsequently, we derive a microscopic model where phase transitions between different locally ordered phase can take place. Here, by performing a numerical analysis we explicitly derive, among the others, the gaps, local order parameters and correlation functions behaviors which allow unambiguously proving the presence of partially gapped deconfined quantum critical points. Our results thus provide new interesting insights on the widely investigated topic of quantum phase transitions.
@@ I Introduction
The mechanism by which one physical state is transformed into a different one, namely a phase transition, represents one of the most fundamental concepts in several areas of physics  Domb et al. (1972); Sachdev (2011). In many cases, phase transitions are realized when an order parameter capturing the properties of a physical state abruptly vanishes, thus signaling the appearance of a phase with new features. This specific mechanism is usually defined as a first order or, equivalently, a discontinuous phase transition  Binder (1987). Notably, a large variety of phase transitions displaying different behaviors exist. Paradigmatic examples are for instance the Berezinskii–Kosterlitz–Thouless  Berezinsky (1971); Kosterlitz and Thouless (1973), Ising  Landau (1937) and, the more exotic, symmetric mass generation  Wang and You (2022) phase transitions.
@@
Within this diversity, the Landau–Ginzburg-Wilson symmetry-breaking paradigm Landau et al. (1999); Wilson and Kogut (1974) rigorously predicts the appearance of discontinuous phase transitions connecting locally ordered (LO) phases with different symmetry properties. Nevertheless, pioneering analysis proved Senthil et al. (2004a, b) that it is possible to go beyond this paradigm. In particular, quantum fluctuations can generate second order continuous phase transitions between distinct gapped LO phases. This implies the presence of a gap closing, or equivalently, the simultaneous and continuous vanishing of the two order parameters characterizing the different LO phases, at a single point: the deconfined quantum critical point (DQCP) Senthil (2023). As proved for two dimensional systems Senthil et al. (2004a, b), here very fascinating and intriguing phenomena including charge fractionalization and emergent gauge fields can appear. In addition, DQCPs are further characterized by the phenomena of symmetry enrichment, meaning that they exhibit symmetry properties not present in any of the two LO phases they connect. This extraordinary variety of physical effects generated an intense theoretical effort which unveiled DQCPs in two-dimensional spin Sandvik (2007); Jiang et al. (2008); Lou et al. (2009); Banerjee et al. (2010); Sandvik (2010); Harada et al. (2013); Chen et al. (2013); Nahum et al. (2015); Shao et al. (2016); Wang et al. (2017); Lee et al. (2019); Wang et al. (2022); Song et al. (2023) and fermionic Li et al. (2019); Assaad and Grover (2016); Liu et al. (2023); Liao et al. (2023) models as well as in 3D Charrier et al. (2008); Sreejith and Powell (2015), and 1D Jiang and Motrunich (2019); Roberts et al. (2019); Huang et al. (2019); Mudry et al. (2019); Roberts et al. (2021); Lee et al. (2023); Baldelli et al. (2024); Romen et al. (2024) Hamiltonians. Notably, these previous studies all focused on systems where only a gap dictates the properties of phases with different symmetries. As a consequence, the gap closing at the transition point was always associated to fully gapless DQCPs where all the correlations have an algebraic decay. This is completely justified by the fact that in the large majority of possible LO phases, the gaps capturing different degrees of freedom are always intrinsically related.
In this regard, fermionic systems deserve special attention. In one dimension, interacting fermions can display the celebrated phenomenon of spin-charge separation  Tomonaga (1950); Luttinger (1963); Haldane (1981). This accounts for the decoupling of spin and charge degrees of freedom which can therefore behave completely independently one from the other. It then appears fundamental to understand whether DQCPs can exist in presence of spin-charge separation and how this phenomenon might affect their properties.
@@
Here, we tackle this fundamental problem. Crucially, we find that spin-charge separation makes possible the appearance of a novel kind of DQCPs which, contrary to the previously known examples, have the special feature of being partially gapped. We prove this by considering in section II the sine-Gordon model which well describes generic interacting one dimensional fermions in the low energy limit. By relying on a field theory analysis, also called bosonization  Alexander O. Gogolin (1998); Giamarchi (2004), we discuss how for conserved total particle number and spin magnetization, three possible fully gapped LO phases can take place: a charge density wave (CDW), an antiferromagnet (AF) and a bond-order-wave (BOW) insulator. Here, bosonization is able to prove the following key aspects: a) power law decay of specific correlation functions can occur at the phase transitions between the three different LO phases, thus proving the presence of at least one gap closing in one single point; b) in two of these transition points, one gap closes while the other one remains finite as signaled by the long-range behaviour of one nonlocal parity correlator; c) at each of the three transition points one or two additional continuous symmetries emerge.
In section III we further enforce our findings by deriving a microscopic model where the three LO phases can take place. Here, in addition to bosonization, we perform a Variational Uniform Matrix Product States (VUMPS)  Haegeman et al. (2013); Zauner-Stauber et al. (2018) based numerical investigation. This method has the advantage that it can mimic the thermodynamic limit allowing for a systematic scaling procedure close to phase transitions Tagliacozzo et al. (2008); Rams et al. (2018); Vanhecke et al. (2019). Notably, this analysis rules out a possible second order phase transition where both the charge and the spin gap vanish in one point. Here, we indeed detect a discontinuous phase transition fully consistent with the Landau’s predictions. At the same time, our results show that continuous phase transitions where one gap closes and the other remains finite characterize our model. On one hand, this result proves the appearance of an emergent $SU(2)$ symmetry in the channel associated with the gap closing. On the other, the presence of an additional finite gap induces long-range order of a parity correlator. Moreover, by performing a bond-dimension scaling extrapolation, we accurately extract the critical exponents which govern how the local order parameters vanish at the transition point. Thanks to this analysis, we find values not compatible with the Ising universality class thus unambiguously proving the appearance of one dimensional partially gapped DCQPs.
@@
Disordered Phases Symmetry Protected Topological Phases Locally Ordered Phases
Luther-Emery Liquid Mott Insulator Haldane Insulator Haldane Liquid Spin Haldane Insulator Charge Density Wave Antiferromagnet Bond Order Wave
Order Parameter nonlocal $\mathcal{O}_{P}^{s}$ nonlocal $\mathcal{O}_{P}^{c}$ nonlocal $\mathcal{O}_{S}^{c}$ nonlocal $\mathcal{O}_{S}^{s}$ nonlocal $\mathcal{O}_{S}^{c}$, $\mathcal{O}_{S}^{s}$ local $\mathcal{O}_{CDW}$ local $\mathcal{O}_{AF}$ local $\mathcal{O}_{BOW}$
$\varphi_{c}$ unpinned 0 unpinned $\sqrt{\pi/8}$ $\sqrt{\pi/8}$ 0 $\sqrt{\pi/8}$ 0
$\varphi_{s}$ 0 unpinned $\sqrt{\pi/8}$ unpinned $\sqrt{\pi/8}$ 0 0 $\sqrt{\pi/8}$
$\Delta_{c}$ $=0$ $\neq 0$ $\neq 0$ $=0$ $\neq 0$ $\neq 0$ $\neq 0$ $\neq 0$
$\Delta_{s}$ $\neq 0$ $=0$ $=0$ $\neq 0$ $\neq 0$ $\neq 0$ $\neq 0$ $\neq 0$
Table 1: List of all the possible gapped phase in one-dimensional fermionic systems with $U(1)$ charge and spin symmetry described by the Hamiltonian in Eq.(1) and derived by studying the microscopic model in Eq.(16). These phases are distinguished by considering the presence of different gaps and their order parameters.
@@ II.1 Derivation of phases with nonlocal or local order
Interacting fermionic systems allow for a field theory description, also called bosonization  Alexander O. Gogolin (1998); Giamarchi (2004), which results accurate in the low energy limit. In particular, a generic lattice microscopic Hamiltonian conserving the total magnetization and particles number can be rewritten in the continuum limit as two decoupled sine-Gordon models ${\cal H}=\sum_{\nu=c,s}{\cal H}_{\nu}^{SG}$ capturing the spin ($s$) and charge ($c$) degrees of freedom
@@
see Appendix A for more details. Here, $v_{\nu}$, $K_{\nu}$ and $g_{\nu}$ are the excitation velocities, Luttinger parameters, and coupling amplitudes respectively, which depend on the microscopic Hamiltonian parameters. At the same time, the behavior of the bosonic fields $\varphi_{\nu}(x)$ and $\vartheta_{\nu}(x)$ governs the appearance of gapped phases, see A . Specifically, a finite gap $\Delta_{\nu}$ in the $\nu$ channel develops when $\varphi_{\nu}(x)$ is pinned to a specific value. Notice that in such a case the effective symmetry relative to the $\nu$ channel is $U(1)$. On the other hand, if $\varphi_{\nu}(x)$ remains unpinned, the $\nu$ degrees of freedom are gapless and an emergent $U(1)\times U(1)\cong SU(2)$  111notice that the effective $SU(2)$ symmetry holds in the case of $SU(2)$ symmetric choices of the microscopic parameters of a Hamiltonian symmetry appears. In this regard, eq. (1) provides different solutions describing gapped phases.
We start by discussing the following cases: a) $\varphi_{c}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{s}$ is unpinned; b) $\varphi_{s}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{c}$ is unpinned; c) $\varphi_{c}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{s}=\sqrt{\frac{\pi}{8}}$. It turns out that gapped phases captured by such specific value $\sqrt{\frac{\pi}{8}}$ of the pinned field $\varphi_{\nu}(x)$ can be uniquely identified through the nonlocal string operator
@@
in the microscopic and bosonized version respectively Montorsi and Roncaglia (2012); Barbiero et al. (2013), where $S_{j}^{s}=(n_{j\uparrow}-n_{j\downarrow})$, $S_{j}^{c}=(1-n_{j})$ and $n_{j}=n_{j,\uparrow}+n_{j,\downarrow}$ are fermionic number operators. Specifically, states of matter which ordering is captured solely by a finite string order parameter
[…]
can be recognized  Montorsi et al. (2017) as symmetry protected topological (SPT) phases  Pollmann et al. (2012); Tang and Wen (2012); Wen (2014); Gu and Wen (2014); Senthil (2015); Shapourian et al. (2017). Based on the formal analogy with the SPT phases occurring in spin-$1$ Heisenberg  Haldane (1983); Affleck et al. (1987); Barbiero et al. (2017) and Bose-Hubbard models  Dalla Torre et al. (2006); Rossini and Fazio (2012); Sugimoto et al. (2019); Fraxanet et al. (2022), the a) solution thus describes a charge gapped Haldane insulator. A spin gapped SPT regime also known as Haldane liquid  Fazzini et al. (2019); Montorsi et al. (2020) emerges when only $\varphi_{s}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{c}$ keeps fluctuating, i. e. b). Finally, c) implies that both $\mathcal{O}_{S}^{\nu}\neq 0$ and therefore a fully gapped spin Haldane insulator phase analogous to the topological regime of spinful SSH models  Anfuso and Rosch (2007); Manmana et al. (2012); Barbiero et al. (2018); Julià-Farré et al. (2022) takes place. Notably, this intrinsic nonlocality makes SPT phases the paradigmatic example of states escaping the Landau’s theory of LO states of matter.
Two additional phases characterized by the absence of local order, and therefore also not captured by the Landau’s theory, occur when $\varphi_{c(s)}=0$ and $\varphi_{s(c)}$ is unpinned. Upon defining the nonlocal parity, also called in higher dimension disorder  Kadanoff and Ceva (1971); Nussinov and Ortiz (2009); Fradkin (2017), operators
@@
[…]
represents the order parameters of such phases. More in detail, these two solutions describe: a spin gapped Luther-Emery liquid (LEL) phase where $\varphi_{s}=0$ and $\varphi_{c}$ is unpinned so that only $\mathcal{O}_{P}^{s}\neq 0$; a Mott insulator (MI) where the charge and spin gaps are $\Delta_{c}\neq 0$, $\Delta_{s}=0$, thus reflecting the unpinning of $\varphi_{s}$ and the pinning of $\varphi_{c}$ to zero so that only $\mathcal{O}_{P}^{c}\neq 0$. Because of absence of any local order, we name both LE and MI as disordered (DIS) phases.
@@
[…]
@@
[…]
@@
[…]

### v4

@@ Abstract
Deconfined quantum critical points are intriguing transition points not predicted by the Landau–Ginzburg-Wilson symmetry-breaking paradigm which are usually identified by the appearance of a continuous phase transition between locally ordered phases. Here, we reveal the presence of deconfined quantum critical points with unexplored properties. Contrary to previously known examples, we show that the phenomenon of spin-charge separation peculiar to interacting low dimensional fermions can allow for the appearance of partially gapped deconfined quantum critical points. We first infer this point by performing a field theory analysis of generic one-dimensional fermionic systems in the low energy limit. Subsequently, we derive a microscopic model where phase transitions between different locally ordered phases can take place. Here, by performing a numerical analysis we explicitly derive, among others, the gaps, local order parameters and correlation functions behavior, supporting the presence of partially gapped deconfined quantum critical points. Our results thus provide new interesting insights on the widely investigated topic of quantum phase transitions.
@@ I Introduction
The mechanism by which one physical state is transformed into a different one, namely a phase transition, represents one of the most fundamental concepts in several areas of physics  [1, 2]. In many cases, phase transitions are realized when an order parameter capturing the properties of a physical state abruptly vanishes, thus signaling the appearance of a phase with new features. This specific mechanism is usually defined as a first order or, equivalently, a discontinuous phase transition  [3]. Notably, a large variety of phase transitions displaying different behaviors exist. Paradigmatic examples are for instance the Berezinskii–Kosterlitz–Thouless  [4, 5], Ising  [6] and, the more exotic, symmetric mass generation  [7] phase transitions.
@@
Within this diversity, the Landau–Ginzburg-Wilson symmetry-breaking paradigm [8, 9] rigorously predicts the appearance of discontinuous phase transitions connecting locally ordered (LO) phases with different symmetry properties. However, pioneering analysis proved [10, 11] that it is possible to go beyond this paradigm. In particular, quantum fluctuations can generate second order continuous phase transitions between distinct gapped LO phases. This implies the presence of a gap closing, or equivalently, the simultaneous and continuous vanishing of the two order parameters characterizing the different LO phases, at a single point: the deconfined quantum critical point (DQCP) [12]. As proved for two-dimensional systems [10, 11], here very fascinating phenomena including charge fractionalization and emergent gauge fields can appear. In addition, DQCPs are further characterized by the phenomena of symmetry enrichment, meaning that they exhibit symmetry properties not present in any of the two LO phases they connect. This extraordinary variety of physical effects generated an intense theoretical effort which unveiled DQCPs in two-dimensional spin [13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25] and fermionic [26, 27, 28, 29] models, as well as in 3D [30, 31, 32], and 1D [33, 34, 35, 32, 36, 37, 38, 39, 40] Hamiltonians. Notably, these previous studies all focused on systems where only a gap dictates the properties of phases with different symmetries. As a consequence, the gap closing at the transition point was always associated with fully gapless DQCPs, where all the correlations have algebraic decay. This is completely justified by the fact that, in the large majority of possible LO phases, the gaps capturing different degrees of freedom are always intrinsically related. Importantly, recent analysis [41, 42] based on conformal bootstrap [43] challenged the existence of two dimensional DQCPs, while the same technique seems to confirm their presence in 3D [44]. These analysis thus motivate even more the characterization of DQCPs in 1D where highly accurate analytical and numerical techniques are available. Moreover, exact results have been recently produced [38]. In such a context, fermionic systems deserve special attention. Indeed, many-body fermionic Hamiltonians can host the celebrated phenomenon of spin-charge separation  [45, 46, 47]. This accounts for the decoupling of spin and charge degrees of freedom which can therefore behave completely independently one from the other. Considering all these aspects, it becomes crucial to understand whether DQCPs can exist in the presence of spin-charge separation, and how this phenomenon may influence their properties.
@@
Here, we tackle this fundamental problem. Crucially, we find that spin-charge separation makes possible the appearance of a novel kind of DQCPs which, contrary to the previously known examples, have the special feature of being partially gapped. We prove this by considering in section II the sine-Gordon model which well describes generic interacting one dimensional fermions in the low energy limit. By relying on a bosonization analysis  [48, 49], we discuss how for conserved total particle number and spin magnetization, three possible fully gapped LO phases can take place: a charge density wave (CDW), an antiferromagnet (AF) and a bond-order-wave (BOW) insulator. Here, bosonization is able to prove the following key aspects: a) power law decay of specific correlation functions can occur at the phase transitions between the three different LO phases, thus proving the presence of at least one gap closing in one single point; b) in two of these transition points, one gap closes while the other one remains finite as signaled by the long-range behavior of one nonlocal parity correlator; c) at each of the three transition points one or two additional continuous symmetries emerge.
In section III we further enforce our findings by deriving a microscopic model where the three LO phases can take place. Here, in addition to bosonization, we perform a Variational Uniform Matrix Product States (VUMPS)  [50, 51] based numerical investigation. This method has the advantage that it can directly simulate systems in the thermodynamic limit, allowing for a systematic scaling procedure close to phase transitions [52, 53, 54]. In this way, we unveil DQCPs where the concurrent disappearance of two local order parameters are related to the closing of a single gap, either the spin or charge one. Moreover, this analysis rules out a possible second order phase transition where both gaps vanish in one point. In this latter scenario, we indeed observe a first order phase transition as predicted by the Landau theory discuss the previous sentence. On one hand, this result proves the appearance of an emergent $U(1)\times U(1)$ symmetry in the channel associated with the gap closing. On the other, the presence of an additional finite gap induces long-range order of a parity correlator. Moreover, by performing a bond-dimension scaling extrapolation, we accurately extract the critical exponents which govern how the local order parameters vanish at the transition point. Thanks to this analysis, we find values not compatible with the Ising universality class thus supporting the appearance of one dimensional partially gapped DCQPs.
@@
Symmetry Protected Topological Phases Disordered Phases Locally Ordered Phases
Haldane
Insulator
Haldane
Liquid
Spin Haldane
Insulator
Luther-Emery
Liquid
Mott
Insulator
Charge Density
Wave
Antiferromagnet
Bond Order
Wave
Order
Parameter
nonlocal
$\mathcal{O}_{S}^{c}$
nonlocal
$\mathcal{O}_{S}^{s}$
nonlocal
$\mathcal{O}_{S}^{c}$, $\mathcal{O}_{S}^{s}$
nonlocal
$\mathcal{O}_{P}^{s}$
nonlocal
$\mathcal{O}_{P}^{c}$
local
$\mathcal{O}_{CDW}$
local
$\mathcal{O}_{AF}$
local
$\mathcal{O}_{BOW}$
$\varphi_{c}$ $\sqrt{\pi/8}$ unpinned $\sqrt{\pi/8}$ unpinned 0 $\sqrt{\pi/8}$ 0 0
$\varphi_{s}$ unpinned $\sqrt{\pi/8}$ $\sqrt{\pi/8}$ 0 unpinned 0 $\sqrt{\pi/8}$ 0
$\Delta_{c}$ $\neq 0$ 0 $\neq 0$ 0 $\neq 0$ $\neq 0$ $\neq 0$ $\neq 0$
$\Delta_{s}$ 0 $\neq 0$ $\neq 0$ $\neq 0$ 0 $\neq 0$ $\neq 0$ $\neq 0$
Table 1: List of all the possible gapped phases in one-dimensional fermionic systems with $U(1)$ charge and spin symmetry described by the Hamiltonian in Eq.(1) and derived by studying the microscopic model in Eq.(17). These phases are distinguished by considering the presence of different gaps and their order parameters.
@@ II.1 Derivation of phases with nonlocal or local order
Interacting fermionic systems allow for a field theory description, also called bosonization  [48, 49], accurate in the low energy limit. In particular, a generic lattice microscopic Hamiltonian conserving the total magnetization and particles number can be rewritten in the continuum limit as two decoupled sine-Gordon models ${\cal H}=\sum_{\nu=c,s}{\cal H}_{\nu}^{SG}$ capturing the spin ($s$) and charge ($c$) degrees of freedom
@@
see Appendix A for more details. Here, $v_{\nu}$, $K_{\nu}$ and $g_{\nu}$ are the excitation velocities, Luttinger parameters, and coupling amplitudes respectively, which depend on the microscopic Hamiltonian parameters. At the same time, the behavior of the bosonic fields $\varphi_{\nu}(x)$ and $\vartheta_{\nu}(x)$ governs the appearance of gapped phases, see A . Specifically, a finite gap $\Delta_{\nu}$ in the $\nu$ channel develops when $\varphi_{\nu}(x)$ is pinned to a specific value. Notice that in such a case the effective symmetry relative to the $\nu$ channel is $U(1)$ (see Appendix). On the other hand, if $\varphi_{\nu}(x)$ remains unpinned, the $\nu$ channel is gapless and an emergent $U(1)\times U(1)$  111notice that this may be enhanced to a full $SU(2)$ symmetry in case of symmetric choices of the microscopic parameters of a Hamiltonian[32] symmetry appears. In this regard, eq. (1) provides different solutions describing gapped phases.
We start by discussing the following cases: a) $\varphi_{c}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{s}$ is unpinned; b) $\varphi_{s}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{c}$ is unpinned  [^2] Importantly, this scenario has also been described as a gapless SPT regime [115, 116]; c) $\varphi_{c}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{s}=\sqrt{\frac{\pi}{8}}$. It turns out that gapped phases captured by such specific value $\sqrt{\frac{\pi}{8}}$ of the pinned field $\varphi_{\nu}(x)$ can be uniquely identified through the nonlocal string operator
@@
in the microscopic and bosonized version respectively [57, 58], where
$S_{j}^{s}=n_{j\uparrow}-n_{j\downarrow},\quad S_{j}^{c}=1-n_{j},$ (3)
are fermionic number operators. Here, $n_{j}=n_{j,\uparrow}+n_{j,\downarrow}$ is the total number of particles. Specifically, states of matter which ordering is captured solely by a finite string order parameter
$\mathcal{O}_{S}^{\nu}=\frac{1}{L}\sum_{j}\langle O_{S}^{\nu}(j)\rangle\rightarrow\mathcal{O}_{S}^{\nu}\sim\frac{1}{L}\int dxO_{S}^{\nu}(x)$ (4)
can be recognized  [59] as symmetry protected topological (SPT) phases  [60, 61, 62, 63, 64, 65]. Based on the formal analogy with the SPT phases occurring in spin-$1$ Heisenberg  [66, 67, 68] and Bose-Hubbard models  [69, 70, 71, 72], the a) solution thus describes a charge gapped Haldane insulator. A spin gapped SPT regime also known as Haldane liquid  [73, 74] emerges when only $\varphi_{s}=\sqrt{\frac{\pi}{8}}$ and $\varphi_{c}$ keeps fluctuating, i. e. b). Finally, c) implies that both $\mathcal{O}_{S}^{\nu}\neq 0$ and therefore a fully gapped spin Haldane insulator phase analogous to the topological regime of spinful SSH models  [75, 76, 77, 78] takes place. Notably, this intrinsic nonlocality makes SPT phases the paradigmatic example of states escaping the Landau’s theory of LO states of matter.
Two additional phases characterized by the absence of local order, and therefore also not captured by the Landau’s theory, occur when $\varphi_{c(s)}=0$ and $\varphi_{s(c)}$ is unpinned. Upon defining the nonlocal parity, also called in higher dimension disorder  [79, 80, 81], operators
@@
$\mathcal{O}^{\nu}_{P}=\frac{1}{L}\sum_{j}\langle O_{P}^{\nu}(j)\rangle\rightarrow\mathcal{O}^{\nu}_{P}\sim\frac{1}{L}\int dxO_{P}^{\nu}(x)$ (6)
represents the order parameters of such phases. More in detail, these two solutions describe: a spin gapped Luther-Emery liquid (LEL) phase where $\varphi_{s}=0$ and $\varphi_{c}$ is unpinned so that only $\mathcal{O}_{P}^{s}\neq 0$; a Mott insulator (MI) where the charge and spin gaps are $\Delta_{c}\neq 0$, $\Delta_{s}=0$, thus reflecting the unpinning of $\varphi_{s}$ and the pinning of $\varphi_{c}$ to zero so that only $\mathcal{O}_{P}^{c}\neq 0$. Because of absence of any local order, we name both LE and MI as disordered phases.
@@
$\mathcal{O}_{CDW}$ $\sim$ $\frac{1}{L}\int dx\langle\sin(\sqrt{2\pi}\varphi_{c}(x))\cos(\sqrt{2\pi}\varphi_{s}(x))\rangle.$
@@
$\mathcal{O}_{AF}$ $\sim$ $\frac{1}{L}\int dx\langle\cos(\sqrt{2\pi}\varphi_{c}(x))\,\sin(\sqrt{2\pi}\varphi_{s}(x))\rangle$
@@
$\mathcal{O}_{BOW}$ $\sim$ $\frac{1}{L}\int dx\langle\cos(\sqrt{2\pi}\varphi_{c}(x))\,\cos(\sqrt{2\pi}\varphi_{s}(x))\rangle$